A 2.0m length of metal wire is connected to a 1.5 V battery, and a current of 5.0mA flows through it. What is the diameter of the wire? (The resistivity of metal is 2.24
Answer #1

R = V/I = 1.5/ 0.005 = 300 ohm

R = resisitivity *( L/ (pi *r2))

r=6.89 micro meter

diameter = 2r =1.38 *10^-5 m
Answer #2

V = I * r *L/pi * r^2

r = 6.89 * 10 ^2

d = 13.79 * 10^2
Answer #3

Resistance = resistiviy x length / area

resistance by equation = 1.5 / 5 x 10-3 = 3 x 102 ohm

3 x 102 = 2.24 x 108 x 2 / pi x r2

Solving we get r = 6.89 x 103 m

ANS diameter = 1.38 x 104 m
Answer #4

R = V/I

R = (1.5V)/(5mA) = 300 Ohms

1 P - resistivity -Length A -cross sectional area

A = pL/R = (2.24*108)(2)/300 = 1.6128

Area of a circle = d2(pi)/4

d = sqrt(4*1.6128/pi) = 1.433m